Smartphones: A poll agency reports that 32% of teenagers aged 12-17 own smartphones. A random sample of 100 teenagers is drawn. Round your answers to at least four decimal places as needed. (e) Find the probability that less than 35% of sampled teenagers own smartphones. The probability that less than 35% of sampled teenagers own smartphones is?
Answer #1

Here, μ = 0.32, σ = 0.0466 and x = 0.35. We need to compute P(X <= 0.35). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z = (0.35 - 0.32)/0.0466 = 0.64

Therefore,
P(X <= 0.35) = P(z <= (0.35 - 0.32)/0.0466)
= P(z <= 0.64)
= 0.7389